You are dealt a single card from a 52 deck of cards. Find the probability that you are dealt a red or a ten card.
Answer by fractalier(6550)

P(red or ten) = P(red) + P(ten) - P(red)*P(ten)
= 1/2 + 1/13 - 1/26 = 26/52 + 4/52 - 2/52 = 28/52 = 7/13